OpenStudy (anonymous):

Which is a set of roots for the quadratic equation x^2+10x-24=0

OpenStudy (anonymous):

can you factor the left side?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

so what two numbers multiply to -24 and add to 10?

OpenStudy (anonymous):

-24 = ??? * ???? 10 = ??? + ????

OpenStudy (anonymous):

-12*2=-24

OpenStudy (anonymous):

5+5=10

OpenStudy (anonymous):

12 * -2 = -24 12 + -2 = 10 they have to be the same numbers for both...

OpenStudy (anonymous):

6 and 4

OpenStudy (anonymous):

it takes practice but that's the two numbers you're looking for 12 and -2.

OpenStudy (anonymous):

ohh thanks :D

OpenStudy (anonymous):

so the equation: x^2+10x-24=0 can be factored as x^2+10x-24=0 (x+12)(x-2) = 0 now it's easy to solve... x+12 = 0 and x-2=0 solve these two to get the solution for your original question.... :)

OpenStudy (anonymous):

you ok from here? what's the solution set?

OpenStudy (anonymous):

{12,-2}

OpenStudy (anonymous):

^was the answer

OpenStudy (anonymous):

no... it's {-12, 2}

OpenStudy (anonymous):

ohh..

OpenStudy (anonymous):

you have to solve that last two equations: x+12=0 and x-2=0

OpenStudy (anonymous):

that's why it's -12 and 2 and not the other way around... :)

OpenStudy (anonymous):

ohh ok.. thanks for the help

OpenStudy (anonymous):

yw...:)